OpenStudy (anonymous):

The average value of a certain type of automobile was $13,140 in 1992 and depreciated to $9420 in 1997. Let y be the average value of the automobile in the year x, where x = 0 represents 1992. Write a linear equation that models the value of the automobile in terms of the year x. A) y = -744x + 9420 B) y = - 1/744x - 9420 C) y = -744x + 13,140 D) y = -744x + 5700

OpenStudy (anonymous):

anything?

OpenStudy (darkprince14):

try to test each given equation with x =5...it's kinda the last resort coz i forgot how to solve problems like this..sorry..

OpenStudy (anonymous):

@jim_thompson5910 can you help?

OpenStudy (anonymous):

do you know how to find the slope of a line if I give you two points?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

when x = 0, y = 13140

jimthompson5910 (jim_thompson5910):

so plugging in x = 0 into choice A for example gives us y = -744x + 9420 y = -744(0) + 9420 y = 0 + 9420 y = 9420 which is NOT the answer (we want y = 13140)

jimthompson5910 (jim_thompson5910):

so choice A is out

OpenStudy (anonymous):

oh

OpenStudy (darkprince14):

that's what they call trial and error XD..

OpenStudy (anonymous):

so is it c.

OpenStudy (darkprince14):

try the given equation on letter be and see if it works... do the same for c and d...

jimthompson5910 (jim_thompson5910):

it is C

OpenStudy (anonymous):

yay! thank you guys!

hero (hero):

Hottmess!, you need to go back to that question you think you finished, but didn't.